OpenStudy (anonymous):

(x-16)+5=-15 What is the solution to the equation?

OpenStudy (anonymous):

(x-16)+5= -15 x-16= -15-5 x-16 = -20 x= -20+16 x= -4

OpenStudy (mathstudent55):

Is this the equation? \((x - 16) + 5 = -15\)

OpenStudy (anonymous):

yes @mathstudent55

OpenStudy (anonymous):

And thank you @PrincessHush

OpenStudy (anonymous):

You are quite welcome(:

OpenStudy (mathstudent55):

First, the parentheses on the left side are not needed, so we just take them off. \(x - 16 + 5 = -15\) Now we combine like terms (-16 and 5) on the left side. \(x - 11 = -15\) Add 11 to both sides: \(x = -4\)